OpenStudy (anonymous):

simplify the radical expression sqrt21y x 5sqrt49y ?

OpenStudy (anonymous):

the x is meant to be a multiplication sign

OpenStudy (kc_kennylau):

\[\sqrt{21y}*5\sqrt{49y}=5\sqrt{21y*49y}=5y*7\sqrt{21}=35y\sqrt{21}\]

OpenStudy (anonymous):

want to help with more?:)